OpenStudy (anonymous):

Rewrite the number in scientific notation: 1042 10³ i could of sworn thats right but its not

OpenStudy (phi):

what number do you want in sci notation?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

idk thats why i got it wrong lol

OpenStudy (anonymous):

i think its 1.04 × 10^3

OpenStudy (phi):

if the question is rewrite 1042*10^3 you want a number between 1 and 10 out front. the way to do that is move the decimal point of 1042.0 left 3 places. if you make the number out front smaller, make the exponent on the 10 bigger by the same number of spaces you move the decimal point

OpenStudy (anonymous):

no its just rewrite 1042

OpenStudy (phi):

in that case, think of 1042 as 1042*10^0 use the same rule: move the decimal point to make the number between 1 and (less than) 10 count the number of spaces you move, and add that number to the exponent of 10 (which is 0 here) (10^0 is the same a 1)

OpenStudy (anonymous):

ok so whats the answer i keep getting10³

OpenStudy (phi):

what is the number out front?

OpenStudy (phi):

you know that 10^3 means 10*10*10 = 1000 1000 is not 1042 so you know that is not correct you need to put a number out front of 10^3 so that you get 1042 when you multiply by 10^3 the number you put out front is 1042 with the decimal point shifted

OpenStudy (anonymous):

1?

OpenStudy (phi):

you are making a mountain out of a molehill think about this: 1234 = 1234*1 1234= 123.4*10 (try it) 1234= 12.34*100 (try it) 1234= 1.234*1000 (try it) the last has a number out front that is between 1 and (less than) 10 so that is scientific notation. but because people don't like writing lots of zeros, they write it as 1.234*10^3 now to 1042

OpenStudy (anonymous):

thankkkkk yyou :D